# Towards Foundation Models for 3D Vision: How Close Are We?

### v1

@@ 4.5 Keypoint Matching
  • Keypoint Detection We investigate what kinds of points humans are more likely to choose when asked to find matching points between two images. We find that humans have a tendency towards choosing salient points. Fig. 6 shows that the subjects were much more likely to choose corners, SIFT keypoints, FAST keypoints, or SuperPoint keypoints than random choice. Among the detectors, SuperPoint has the best alignment with human keypoint selection. It shows that neural networks have higher similarities to humans than hand-crafted detectors.
@@ 5 Limitations
There are a few limitations of our analysis which we plan to address in future work: 1) While we focus on the four tasks that we consider most fundamental for 3D vision—relative depth estimation, spatial VQA, relative camera pose estimation, and keypoint matching—it would be beneficial to evaluate additional tasks, such as surface normals. 2) Due to time and resource constraints, we only evaluated closed-source VLMs. Evaluating open-source VLMs could provide a more comprehensive comparison.

### v2

@@ 4.5 Keypoint Matching
  • Keypoint Detection We investigate what kinds of points humans are more likely to choose when asked to find matching points between two images. We find that humans have a tendency towards choosing salient points. Fig. 6 shows that the subjects were much more likely to choose corners, SIFT keypoints, FAST keypoints, or SuperPoint keypoints than random choice. Among the detectors, SuperPoint has the best alignment with human keypoint selection. This shows that neural networks have higher similarity to humans than hand-crafted detectors.
@@ 5 Limitations
There are a few limitations in our analysis: 1) While we focus on the four tasks that we consider most fundamental for 3D vision—relative depth, spatial reasoning, relative camera pose estimation, and keypoint matching—it would be beneficial to evaluate additional tasks, such as surface normals. 2) Due to time and resource constraints, we only evaluated closed-source VLMs. Evaluating open-source VLMs could provide a more comprehensive comparison.